OpenStudy (anonymous):

8H2S(g)+4O2(g)→8H2O(g)+S8(g) Δ[H2S]/Δt = -0.044M/s Find Δ[S8]/Δt , I thought I knew the process but I got the wrong answer.. any assistance?

OpenStudy (anonymous):

-1/8 d[H2S]/dt = 1/1 d[S8]/dt =5.5x10^-3

OpenStudy (aaronq):

that should be the right answer.